OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). -5y-x-10 all divide by-5 y=-5+2 am I doing this right? @ agent0smith

OpenStudy (bahrom7893):

well yes, but you divided wrong. What's -x divided by 5?

OpenStudy (anonymous):

pos 5

OpenStudy (bahrom7893):

No.. Let's try something simpler.. what is 2 divided by 5?

OpenStudy (anonymous):

0.4

OpenStudy (bahrom7893):

Write it in fractional form, and don't use a calculator.

OpenStudy (anonymous):

2/5

OpenStudy (bahrom7893):

That's what I was looking for.. How about one divided by nine?

OpenStudy (anonymous):

1/9

OpenStudy (bahrom7893):

Okay, we're getting there.. How about ☺ divided by ☻ ?

OpenStudy (anonymous):

=)

OpenStudy (bahrom7893):

No im kinda serious here

OpenStudy (bahrom7893):

So what is ☺ divided by ☻ ? (You can copy paste the smileys if you can't type them out)

OpenStudy (anonymous):

☺/☻

OpenStudy (bahrom7893):

Yes. Now take a guess... what's -x divided by 5?

OpenStudy (anonymous):

-x/5

OpenStudy (bahrom7893):

Yup. So you had x - 5y = -10. Divide everything by 5. What do you get (you can divide each term individually)?

OpenStudy (anonymous):

-x/5_2

OpenStudy (anonymous):

+2

OpenStudy (bahrom7893):

Equals?

OpenStudy (anonymous):

y

OpenStudy (bahrom7893):

Okay, actually we were off even earlier - I didn't notice that the negative was wrong. So we had: x - 5y = -10 Which turned into (after adding 5y to both sides): x - 5y + 5y = -10 + 5y x = -10 + 5y Which then became (after adding 10 to both sides): x + 10 = -10 +5y + 10 x+10 = 5y And now if you divide everything by 5, you get: x/5 + 2 = y, which is the same as: y = (1/5)x + 2

OpenStudy (bahrom7893):

So what's the slope of that equation?

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

1/5x

OpenStudy (bahrom7893):

It actually is 1/5 A slope of the equation in the form y = Mx+b is equal to M

OpenStudy (bahrom7893):

So we know that M = 1/5

OpenStudy (bahrom7893):

You need to write the equation of the line perpendicular to the line that has slope of 1/5. What do we know about slopes of perpendicular lines?

OpenStudy (anonymous):

y-5=1/5(x-2) ?

OpenStudy (bahrom7893):

close.. that would be correct if you wanted a parallel equation.

OpenStudy (bahrom7893):

i mean an equation for the line parallel to that one. you want it to be perpendicular.

OpenStudy (anonymous):

right, so I am kind of confused on what to do.